# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

### v2

@@ 5. Experiments
  • (RQ1) Can ScaleGNN outperform the state-of-the-art GNN methods regarding predictive accuracy on real-world datasets?
@@
  • (RQ6) How do different hyperparameter settings affect the performance?
5.1. Experimental Settings
5.1.1. Datasets
We have evaluated the effectiveness of ScaleGNN using six real-world graph datasets, including Citeseer, Cora, Pubmed, and ogbn-arxiv, ogbn-products, ogbn-papers100M. The first three datasets (Citeseer, Cora, and Pubmed) are relatively small and have been adopted by existing researches [15], while the latter three (ogbn-arxiv, ogbn-products, and ogbn-papers100M) are large-scale heterogeneous graphs commonly used in scalable HGNN evaluations. The detailed description of these datasets is shown in the Table 2.
5.1.2. Baselines
@@
  • Traditional GNN methods: R-GCN (Schlichtkrull et al., 2018), HetGNN (Zhang et al., 2019), HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), Simple-HGN (Lv et al., 2021), HINormer (Mao et al., 2023).
  • Deep GNN methods: SIGN (Rossi et al., 2020), S2GC (Zhu and Koniusz, 2021), GBP (Chen et al., 2020), GAMLP (Zhang et al., 2022b), GRAND+ (Feng et al., 2022), LazyGNN (Xue et al., 2023).
  • Scalable GNN methods: NARS (Yu et al., 2020), SeHGNN (Yang et al., 2023), RpHGNN (Hu et al., 2024).
Notice that most of traditional and deep GNN methods encounter out-of-memory (OOM) issues when dealing with large datasets (such as products and paper100M). As for R-GCN, while the original version presents OOM issues with large datasets, it is often used as a baseline for large-scale HGNNs. in an effort to adapt R-GCN for large datasets, we follow NARS’ experimental setting and adopt the neighbor sampling strategy used by GraphSAGE (Hamilton et al., 2017). Besides, scalable GNN methods NARS, SeHGNN, and RpHGNN, are both pre-computation-based methods which differ from other end-to-end models. Most of their model runtime is consumed in obtaining the pre-computation tensor for large-scale graphs, with only a tiny amount of time spent on training and inference.
5.1.3. Implementation Details.
Table 3. Performance comparison of all models on six real-world datasets. Mi-F1 and Ma-F1 are short for micro-F1 and macro-F1. Marker * indicates the results are statistically significant against the best-performed baselines (t-test with p-value ¡ 0.01).
@@
RpHGNN 73.12 72.54 83.27 82.37 80.60 79.22 71.98 63.11 80.65 71.34 64.97 56.50
Ours ScaleGNNb 73.11 72.59 83.60 82.41 80.54 79.24 71.95 63.23 80.78 71.34 64.95 56.51
ScaleGNN 74.22* 73.61* 84.53* 84.01* 81.95* 80.50* 73.02* 64.08* 81.64* 72.33* 65.53* 57.29*
5.2. Performance Comparison (RQ1)
We evaluate the performance of our ScaleGNN and all baselines on six real-world datasets under the node classification task. The experimental results are shown in Table 3. The best results are highlighted in bold, and the second-best results are underlined. As we can see, our ScaleGNN achieves the optimal performance, significantly outperforming all baselines in both micro-F1 and macro-F1 metrics on six datasets, with specific emphasis on the improvement relative to the SOTA methods.
Traditional GNN methods work flawlessly on small-scale datasets, but most of them cannot be extended to large-scale graphs and are prone to OOM issues. For example, R-GCN, HAN, Simple-HGN, and HINormer can get experimental results on ogbn-arxiv, but can’t run on larger-scale graphs such as ogbn-products and ogbn-paper100M. For deep GNN methods, they work on learning large-scale graph representations using deeper GNNs while mitigating the over-smoothing problem on large-scale graphs. A common issue with such methods is that they typically convolve dozens or even tens of GNN layers. while capturing the caveat features to some extent, it also greatly increases the computational complexity of the model. Pre-computation-based GNN approaches are distinguished from end-to-end models by the fact that they consume a significant amount of time in pre-computing the tensor of large-scale graphs, so that the models do not need to iteratively learn graph representations during training and inference.
It is worth mentioning that our basic model ScaleGNNb, although retaining only the basic adaptive high-order feature fusion module, achieves very similar performance to one of the SOTA methods, RpHGNN, and runs efficiently in the far outperforms all baselines. ScaleGNN gains further performance gains despite being less efficient than ScaleGNNb. Therefore, ScaleGNN and ScaleGNNb provide an important reference value for us to explore the trade-off between model accuracy and efficiency.
5.3. Efficiency Analysis (RQ2)
Figure 2. Efficiency comparison with SOTA GNN methods, and speedup analysis compared with our ScaleGNNb.
@@ 5.3.1. Computation Efficiency
5.3.2. Contrastive Speedup Analysis
@@
5.4. Mitigating Over-smoothing Issue (RQ3)
@@
Figure 3. Experimental results of our model over SOTA GNN models w.r.t. the number of hops.
@@
The key to ScaleGNN’s resilience against over-smoothing lies in its adaptive feature aggregation mechanism. Unlike conventional approaches that uniformly aggregate neighborhood information, ScaleGNN dynamically balances local and global features, ensuring that useful information is retained while mitigating excessive smoothing. Specifically, our ScaleGNN Selectively incorporates higher-order neighbors with controlled influence, preventing over-mixing of features. It employs adaptive diffusion mechanisms to regulate information propagation, ensuring a balance between short-term and long-term dependencies. ScaleGNN utilizes a local contribution score (LCS) to prioritize informative neighbors, reducing noise and redundancy in feature aggregation. These strategies allow ScaleGNN to maintain a richer set of node representations, even in deep architectures, making it particularly effective for large-scale graphs.
(a) ogbn-products
(b) ogbn-paper100M
(c) ogbn-products
(d) ogbn-paper100M
Figure 4. The parameter sensitivity experimental results.
5.5. Ablation Study (RQ4)
@@
  • w/o Ada removes the adaptive high-order feature fusion module and employs the deep GNN method S2GC (Zhu and Koniusz, 2021) to obtain the graph representation.
  • w/o Low eliminates the low-order enhanced feature aggregation and only employs the adaptive high-order feature fusion.
  • w/o Mask removes the high-order redundant feature masking mechanism.
Table 4. The comparison of ScaleGNN and its variants on micro-F1 and macro-F1 (Mi-F1 and Ma-F1 for short) metrics.
@@
w/o Mask 71.49 62.70 80.42 71.04 64.45 55.98
ScaleGNN 73.02 64.08 81.64 72.33 65.53 57.29
The results demonstrate that each component is vital for improving ScaleGNN’s performance. w/o Ada variant experiences the largest performance drop across all datasets, with reductions of 1.77% in Micro-F1 and 1.53% in Macro-F1 on ogbn-arxiv, 1.16% and 1.54% on ogbn-products, and 1.34% and 1.42% on ogbn-paper100M, highlighting the importance of adaptive high-order feature fusion for capturing structural dependencies and preventing information loss. w/o Low variant also shows notable degradation, especially on ogbn-arxiv and ogbn-paper100M, with Micro-F1 dropping by 0.95% and 0.56%, respectively, indicating that low-order enhanced feature aggregation is crucial for maintaining local information. w/o Mask variant sees moderate degradation, with Micro-F1 decreasing over 1% on ogbn-arxiv, ogbn-products, and ogbn-paper100M, showing the effectiveness of high-order redundant feature masking in reducing redundancy and preventing excessive feature mixing.
Overall, ScaleGNN consistently outperforms all ablated versions, confirming the significant contribution of each module to the model’s success. Adaptive high-order feature fusion is the most influential in improving performance, followed by low-order enhanced feature aggregation and high-order redundant feature masking. The combination of these mechanisms enables ScaleGNN to achieve superior results on various large-scale graph datasets.
@@
(b) ogbn-paper100M
Figure 5. The trade-off between efficiency and accuracy on large-scale graph datasets.
5.6. Trade-off Analysis: Performance vs. Efficiency (RQ5)
In this section, we evaluate the trade-off between efficiency and accuracy for ScaleGNN and ScaleGNNb by analyzing their runtime and Micro-F1 scores across different hop settings.
The runtime analysis reveals that ScaleGNN incurs a higher computational cost than ScaleGNNb across all hop settings. As the number of hops increases, both methods show rising runtimes, reflecting the increased complexity of message passing over larger neighborhoods. However, ScaleGNN consistently requires more computational resources, suggesting that its modeling choices introduce additional overhead.
In terms of performance, ScaleGNN consistently outperforms ScaleGNNb in Micro-F1 scores. Both methods benefit from increasing hops, as it helps capture structural and semantic relationships. However, beyond a certain point, performance gains diminish due to redundant or noisy information from distant nodes.
These results highlight the trade-off between accuracy and efficiency. While ScaleGNN achieves higher Micro-F1 scores, it comes with a significantly higher computational cost. In contrast, ScaleGNNb offers a more efficient alternative with a slight performance compromise. This suggests that model selection should be based on practical constraints: ScaleGNN is ideal for accuracy-focused tasks, while ScaleGNNb is better for resource-constrained environments. Balancing efficiency and accuracy remains crucial in large-scale graph learning.
5.7. Parameter Sensitivity Analysis (RQ6)
5.7.1. The effect of hyperparameters $\beta$
@@
5.7.2. The effect of hyperparameters $\lambda_{1}$ and $\lambda_{2}$
We investigate the sensitivity of $\lambda_{1}$ and $\lambda_{2}$ in the final loss function by varying their values within $\{0.0005,0.001,0.005,0.01,0.05,0.1\}$ and evaluating the Micro-F1 score (%). As shown in Fig. 4, $\lambda_{1}$ is more sensitive, peaking at $\lambda_{1}=0.005$ and $\lambda_{1}=0.01$, with performance declining as $\lambda_{1}$ increases. This suggests that moderate regularization improves performance, while excessive regularization harms it. On the other hand, $\lambda_{2}$ shows a more stable trend, optimizing at $\lambda_{2}=0.01$, with minimal fluctuations, indicating that it primarily enhances model stability. These results highlight the need for careful hyperparameter tuning: $\lambda_{1}$ requires precise adjustment for optimal performance, while $\lambda_{2}$ ensures robustness.
6. Conclusion

### v4

@@ V Experiments
  • (RQ1) Can ScaleGNN outperform SOTA GNN methods regarding predictive accuracy on real-world datasets?
@@
  • (RQ6) How do different hyperparameter settings affect the model performance?
V-A Experimental Settings
V-A1 Datasets
We have evaluated the effectiveness of ScaleGNN using six real-world graph datasets, including Citeseer, Cora, Pubmed, and ogbn-arxiv, ogbn-products, ogbn-papers100M. The first three datasets (Citeseer, Cora, and Pubmed) are relatively small and have been adopted by existing work [15], while the latter three (ogbn-arxiv, ogbn-products, and ogbn-papers100M) are large-scale heterogeneous graphs commonly used in scalable HGNN evaluations. The detailed description of these datasets is shown in Table II.
V-A2 Baselines
@@
Traditional GNN methods:
  • R-GCN [53] introduces relation-specific graph convolution operations to effectively model multi-relational knowledge graphs and achieves strong performance on tasks like link prediction and entity classification.
  • Cluster-GCN [19] is an efficient GCN algorithm that trains models by sampling dense subgraphs partitioned via graph clustering. It significantly reduces memory and computational overhead, enabling efficient training on million-scale graphs for the first time.
  • HetGNN [52] is a heterogeneous graph neural network that jointly learns from both heterogeneous graph structure and node content by aggregating multi-type neighbor and attribute information, achieving superior performance on tasks like link prediction and node classification.
  • HAN [50] employs a hierarchical attention mechanism, combining node-level and semantic-level attention to effectively capture the importance of both meta-path neighbors and meta-paths, enabling expressive node representations in heterogeneous graphs.
  • MAGNN [51] integrates node content, intermediate nodes along meta-paths, and multiple meta-paths through hierarchical aggregation, enabling more expressive and accurate representations in heterogeneous graphs.
  • Simple-HGN [57] is a strong and straightforward baseline for heterogeneous graph learning that, with proper input processing, outperforms previous heterogeneous GNN models across standardized benchmark datasets.
  • HINormer [58] leverages a graph transformer architecture with global-range attention and specialized encoders to effectively capture both structural and heterogeneous relational information, enabling comprehensive node representations in heterogeneous information networks.
Deep GNN methods:
  • SIGN [21] eliminates the need for neighbor sampling by using efficiently precomputed graph convolutional filters of varying sizes, enabling fast and scalable training and inference on large-scale graphs.
  • S2GC [23] uses a modified Markov diffusion kernel to efficiently aggregate information from both local and global neighborhoods, capturing multi-scale context while mitigating over-smoothing in deeper graph convolutions.
  • GBP [24] employs a localized bidirectional propagation strategy to achieve efficient and scalable graph learning, enabling sub-linear time complexity and state-of-the-art performance on massive graphs without relying on sampling.
  • GAMLP [22] adaptively integrates information from multiple neighborhood scales using attention, effectively addressing over-smoothing and enabling highly scalable and efficient graph learning on large-scale and sparse graphs.
  • GRAND+ [71] enhances scalable graph learning by introducing an efficient pre-computation algorithm for propagation and a confidence-aware consistency loss, enabling fast and accurate semi-supervised learning on large-scale graphs.
  • LazyGNN [72] captures long-distance dependencies in graphs using shallow models rather than deeper architectures, effectively avoiding neighborhood explosion and achieving efficient, scalable performance on large-scale graphs.
Scalable GNN methods:
  • NARS [67] efficiently computes neighbor-averaged features over sampled relation subgraphs, enabling scalable and accurate learning on large heterogeneous graphs without complex end-to-end GNN architectures.
  • SeHGNN [68] streamlines heterogeneous graph learning by using pre-computed mean aggregation and a single-layer structure with long meta-paths, combined with transformer-based semantic fusion, achieving high accuracy and fast training with low complexity.
  • RpHGNN [69] combines efficient pre-computation with low information loss by using random projection squashing and a relation-wise even-odd propagation scheme, enabling fast and accurate learning on heterogeneous graphs.
TABLE III: Performance comparison of all models on six datasets. Mi-F1 and Ma-F1 are short for Micro-F1 and Macro-F1. Marker * indicates the results are statistically significant against the best-performed baselines (t-test with p-value $<$ 0.01).
@@
RpHGNN 73.12 72.54 83.27 82.37 80.60 79.32 72.06 63.34 80.89 71.44 64.95 56.50
Ours ScaleGNNb 72.51 72.74 83.42 82.60 80.58 79.25 71.98 63.11 80.75 71.34 64.97 56.60
ScaleGNN 74.44* 73.76* 84.75* 84.21* 82.11* 80.67* 73.25* 64.20* 81.82* 72.47* 65.69* 57.42*
Notice that most of the traditional and deep GNN methods encounter out-of-memory (OOM) issues when dealing with large datasets (such as products and paper100M). As for R-GCN, while the original version presents OOM issues with large datasets, it is often used as a baseline for large-scale HGNNs. To adapt R-GCN for large datasets, we follow NARS’ experimental setting and adopt the neighbor sampling strategy used by GraphSAGE [18]. Besides, scalable GNN methods, NARS, SeHGNN, and RpHGNN, are all pre-computation-based methods that differ from other end-to-end models. Most of their model runtime is consumed in obtaining the pre-computation tensor for large-scale graphs, with only a tiny amount of time spent on training and inference.
V-A3 Implementation Details
V-B Performance Comparison (RQ1)
We evaluate the performance of our ScaleGNN and all baselines on six real-world datasets under the node classification task. The experimental results are shown in Table III. The best results are highlighted in bold, and the second-best results are underlined. As we can see, our ScaleGNN achieves the optimal performance, significantly outperforming all baselines in both micro-F1 and macro-F1 metrics on six datasets, with specific emphasis on the improvement relative to the SOTA methods.
Traditional GNN methods work flawlessly on small-scale datasets, but most of them cannot be extended to large-scale graphs and are prone to OOM issues. For example, R-GCN, HAN, Simple-HGN, and HINormer can get experimental results on ogbn-arxiv, but cannot run on larger-scale graphs such as ogbn-products and ogbn-paper100M. For deep GNN methods, they work on learning large-scale graph representations using deeper GNNs while mitigating the over-smoothing problem on large-scale graphs. A common issue with such methods is that they typically convolve dozens or even tens of GNN layers. While capturing the caveat features to some extent, it also greatly increases the computational complexity of the model. Pre-computation-based GNNs are distinguished from end-to-end models by the fact that they consume a significant amount of time in pre-computing the tensor of large-scale graphs, so that the models do not need to iteratively learn graph representations during training and inference.
It is worth mentioning that our basic model ScaleGNNb, although retaining only the basic adaptive high-order feature fusion module, achieves a very similar performance to the SOTA method RpHGNN and runs efficiently, far outperforming all baselines. ScaleGNN gains further performance despite being less efficient than ScaleGNNb. Therefore, ScaleGNN and ScaleGNNb provide an important reference value for us to explore the trade-off between model accuracy and efficiency.
V-C Efficiency Analysis (RQ2)
Figure 2: Efficiency comparison with SOTA GNN methods, and speedup analysis compared with our ScaleGNNb.
@@ V-C1 Computation Efficiency
V-C2 Contrastive Speedup Analysis
@@
V-D Mitigating Over-smoothing Issue (RQ3)
@@
Figure 3: Experimental results of our model over SOTA GNN models w.r.t. the number of hops.
@@
The key to ScaleGNN’s resilience against over-smoothing lies in its adaptive feature aggregation mechanism. Unlike conventional approaches that uniformly aggregate neighborhood information, ScaleGNN dynamically balances local and global features, ensuring that useful information is retained while mitigating excessive smoothing. Specifically, our ScaleGNN selectively incorporates higher-order neighbors with controlled influence, preventing over-mixing of features. It employs adaptive diffusion mechanisms to regulate information propagation, ensuring a balance between short-term and long-term dependencies. ScaleGNN utilizes LCS to prioritize informative neighbors, reducing noise and redundancy in feature aggregation. These strategies allow ScaleGNN to maintain a richer set of node representations, even in deep architectures, making it particularly effective for large-scale graphs.
V-E Ablation Study (RQ4)
@@
  • w/o Ada removes the adaptive high-order feature fusion module and employs the deep GNN method S2GC [23] to obtain the graph representation.
  • w/o Low excludes the low-order enhanced feature aggregation and only employs adaptive high-order feature fusion.
  • w/o LCS removes the high-order redundant feature masking mechanism.
  • w/o SC removes the masking sparse constraint and keeps the same retained number of nodes in each $\mathbf{A}_{k}^{\text{filter}}$.
TABLE IV: The comparison of ScaleGNN and its variants on micro-F1 and macro-F1 (Mi-F1 and Ma-F1 for short) metrics.
@@
w/o LCS 71.49 62.70 80.42 71.04 64.45 55.98
w/o SC 73.02 64.01 81.58 72.23 65.46 57.19
ScaleGNN 73.25 64.20 81.82 72.47 65.69 57.42
The results demonstrate that each component is vital for improving ScaleGNN’s performance. w/o Ada variant experiences the largest performance drop across all datasets, with reductions of 1.77% in Micro-F1 and 1.53% in Macro-F1 on ogbn-arxiv, 1.16% and 1.54% on ogbn-products, and 1.34% and 1.42% on ogbn-paper100M, highlighting the importance of adaptive high-order feature fusion for capturing structural dependencies and mitigating information loss. w/o Low variant also shows notable degradation, especially on ogbn-arxiv and ogbn-paper100M, with Micro-F1 dropping by 0.95% and 0.56%, respectively, indicating that low-order enhanced feature aggregation is crucial for maintaining local information. w/o LCS variant sees moderate degradation, with Micro-F1 decreasing over three large-scale datasets, showing the effectiveness of high-order redundant feature masking in reducing redundancy and preventing excessive feature mixing. w/o SC, showing that it still falls short of the full-fledged ScaleGNN model despite being the best of all variants, demonstrating the important role of the LCS-based sparse constraints. It employs an attention-based relevance function, which allows ScaleGNN to adaptively assess and filter neighbors in a differentiable and task-driven manner, significantly improving its generalization and scalability in large-scale graphs.
Overall, ScaleGNN consistently outperforms all variants, confirming the significant contribution of each component. Adaptive high-order feature fusion is the most influential in improving performance, followed by low-order enhanced feature aggregation and high-order redundant feature masking. The combination of these mechanisms enables ScaleGNN to achieve superior results on various large-scale graph datasets.
V-F Trade-off Analysis: Performance vs. Efficiency (RQ5)
In this section, we evaluate the trade-off between efficiency and accuracy for ScaleGNN and ScaleGNNb by analyzing their runtime and Micro-F1 scores across diverse hop settings.
@@
(b) ogbn-papers100M
Figure 4: The trade-off between efficiency and accuracy on large-scale graph datasets.
The runtime analysis reveals in Fig. 4 that ScaleGNN incurs a higher computational cost than ScaleGNNb across all hop settings. As the number of hops increases, both methods show rising runtimes, reflecting the increased complexity of message passing over larger neighborhoods. However, ScaleGNN consistently requires more computational resources, suggesting that its modeling choices introduce additional overhead.
ScaleGNN consistently outperforms ScaleGNNb in Micro-F1 scores. Both methods benefit from increasing hops, as it helps capture structural and semantic relationships. However, beyond a certain point, performance gains diminish due to redundant or noisy information from distant nodes. These results highlight the trade-off between accuracy and efficiency. While ScaleGNN achieves higher Micro-F1 scores, it comes with a significantly higher computational cost. In contrast, ScaleGNNb offers a more efficient alternative with a slight performance compromise. This suggests that model selection should be based on practical constraints: ScaleGNN is ideal for accuracy-focused tasks, while ScaleGNNb is better for resource-constrained environments. Balancing efficiency and accuracy remains crucial in large-scale graph learning.
(a) ogbn-products
(b) ogbn-papers100M
(c) ogbn-products
(d) ogbn-papers100M
Figure 5: The parameter sensitivity experimental results.
V-G Parameter Sensitivity Analysis (RQ6)
V-G1 The effect of hyperparameters $\beta$
@@
V-G2 The effect of hyperparameters $\lambda_{1}$ and $\lambda_{2}$
We assess the sensitivity of $\lambda_{1}$ and $\lambda_{2}$ in the loss function by varying their values within {0.0005, 0.001, 0.005, 0.01, 0.05, 0.1} and evaluating Micro-F1. As shown in Fig. 5, ScaleGNN achieves the best performance at $\lambda_{1}=0.005$, with results improving up to this point and declining thereafter, reflecting the benefit of moderate regularization. In contrast, performance with respect to $\lambda_{2}$ is more stable, peaking at $\lambda_{2}=0.01$ with only slight variations, indicating its role in maintaining model stability. These results highlight the need for careful tuning: $\lambda_{1}$ requires precise adjustment for optimal results, while $\lambda_{2}$ ensures robustness. The loss terms weighted by $\lambda_{1}$ and $\lambda_{2}$ complement each other, balancing performance and efficiency.
VI Conclusion